OpenStudy (anonymous):

Can you help me solve this? 2|3d| = 4

OpenStudy (anonymous):

2|3d| = 4 2|3| |d| = 4 6 |d| = 4 |d| = 2/3 d = -2/3 or 2/3

OpenStudy (anonymous):

d=+/- 2/3